OpenStudy (grimsnightmare):

Change mass to number of particles; How many molecules does 36.0 grams of water represent? I know this is a two step problem, 1 - mass to moles, 2 - moles to particles. I am having some problems understanding the information provided to me. is 36.0g the molar mass of H2O? And if it's not I am having problems finding it for hydrogen and oxygen. Please help.

OpenStudy (anonymous):

36.0 gram is not the molar mass. Molar mass of H2O is 18 so mole=mass/molar mass is the formula 36/18 will give you the moles now for particles just multiply the moles by avogadro's number.

OpenStudy (grimsnightmare):

How did you find the 18?

OpenStudy (anonymous):

H2O atomic masses of both elements H and O are given in the periodic Table. O has 16 and H has 1. There are two atoms of H in a compound so 2×1= 2 g. 16 + 2= 18g. Molecular mass is the mass of no. Of all atoms of elements in a compound.

OpenStudy (grimsnightmare):

Alright. Thank you very much for the help.

OpenStudy (anonymous):

:)

OpenStudy (grimsnightmare):

I was just confused because in an example I was provided, we were working on step one problem. C6H12O6 and it was written "If you find the mass of a sample of glucose to be 90.0g, how many moles of glucose is there. She said carbon= 12.0g * 6, hydrogen = 1.01g * 12, Oxygen = 16.6 * 6. then all the final answers put together equaled 180.1g or 180g(1.80 * 10^2 g). Then making the final answer 0.500 moles. Because mass=90.0 and molar mass was 180.

OpenStudy (anonymous):

So now you know what they're saying right whenever they say mass it's not the same thing as molar mass. Whenever i get these type of questions i just work out the molar mass just to be sure.

OpenStudy (anonymous):

If they're going to give you the molar mass of a compound they're sure going to write molar mass. And not just mass/mass of a sample.

OpenStudy (grimsnightmare):

yes and thank you for clearing that up for me. I didn't realize how much the question made sense until just now haha. However I am still confused how my teacher 16.6 for oxygen. I only get 16.0 when rounding.

OpenStudy (anonymous):

Oxygen has 15.999 g in periodic tables 16.6 is wrong because rounding it up will give 17.0. By rounding 15.999 we get 16.0 we do it for simplicity.

OpenStudy (grimsnightmare):

That's what I thought. Thank you again for the help

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

can you have a look at my question.that would be a great help. :)